OpenStudy (anonymous):

if 5 times a number is increased by 5, the result is 30. find the number

mathslover (mathslover):

ok .. so what you tried ?

mathslover (mathslover):

Where r u having problem ?

OpenStudy (anonymous):

yes

OpenStudy (jiteshmeghwal9):

Let the no be 'x'. then, A/q, =>5x+5=30 =>5x=30-5 =>5x=25 =>x=25/5=5.

OpenStudy (jiteshmeghwal9):

clear????

OpenStudy (anonymous):

yes thank you

OpenStudy (jiteshmeghwal9):

yw:)